Mario purchased stock in three different companies. He paid $50 per share for Smith, $45 per share for Jones and $30 per share for Adams, and spent a total of $1020. Mario purchased 2 more shares of Jones than Smith, and twice as many shares of Adams as Smith. How many shares of Jones did he purchase?

Answer by lwsshak3(11628)

let x = number of Jones shares Mario purchased
then,x-2 = number of Smith shares Mario purchased
then,2(x-2) = number Adams shares Mario purchased
$45x+$50(x-2)+$30*2(x-2)=$1020
45x+50x-100+60x-120=1020
155x-220=1020
155x=1240
x=8
ans:Mario purchased 8 shares of Jones stock